Given Density 360 kg/m^3 Diameter 0.7 m e= positive , distence between capacitor 5 mm

Find the potential difference?
Answer #1

diameter, D = 0.7 m

Area of the plates, A = pi*D^2/4

= pi*0.7^2/4

= 0.3848 m^2

Q = +e = 1.6*10^-19 C

d = 5 mm


Capacitnace, C = A*epsilon/d

= 0.3848*8.854*10^-12/(5*10^-3)

= 6.81*10^-10 F

Apply, Q = C*V

==> V = Q/C

= 1.6*10^-19/(6.81*10^-10)

= 2.35*10^-10 volts